In 2018, the Marion Company purchased land containing a mineral mine for $1,640,000. Additional costs of $564,000 were incurred to develop the mine. Geologists estimated that 600,000 tons of ore would be extracted. After the ore is removed, the land will have a resale value of $104,000.

To aid in the extraction, Marion built various structures and small storage buildings on the site at a cost of $252,000. These structures have a useful life of 10 years. The structures cannot be moved after the ore has been removed and will be left at the site. In addition, new equipment costing $91,000 was purchased and installed at the site. Marion does not plan to move the equipment to another site, but estimates that it can be sold at auction for $7,000 after the mining project is completed.

In 2018, 55,000 tons of ore were extracted and sold. In 2019, the estimate of total tons of ore in the mine was revised from 600,000 to 818,000. During 2019, 91,000 tons were extracted.

Required:
1. Compute depletion and depreciation of the mine and the mining facilities and equipment for 2018 and 2019. Marion uses the units-of-production method to determine depreciation on mining facilities and equipment.
2. Compute the book value of the mineral mine, structures, and equipment as of December 31, 2019.
Answer #1

Solution 1&2:

Mine:

Cost to depleted for mine = $1,640,000 + $564,000 - $104,000 = $2,100,000

Depletion expense for 2018 = $2,100,000 * 55000/600000 = $192,500

Book value at the end of 2018 = $2,204,000 - $192,500 = $2,011,500

Remaining cost to be depreciated at beginning of 2019 = $2,100,000 - $192,500 = $1,907,500

Depletion expense for 2019 = $1,907,500 * 91000/(818000-55000) = $227,500

Book value at the end of 2019 = $2,011,500 - $227,500 = $1,784,000

Mining facilities:

Cost to depreciated = $252,000

Depreciation expense for 2018 = $252,000 * 55000/600000 = $23,100

Book value at the end of 2018 = $252,000 - $23,100 = $228,900

Remaining cost to be depreciated at beginning of 2019 = $228,900

Depreciation expense for 2019 = $228,900 * 91000/(818000-55000) = $27,300

Book value at the end of 2019 = $228,900 - $27,300 = $201,600

Mining facilities:

Cost to depreciated = $252,000

Depreciation expense for 2018 = $252,000 * 55000/600000 = $23,100

Book value at the end of 2018 = $252,000 - $23,100 = $228,900

Remaining cost to be depreciated at beginning of 2019 = $228,900

Depreciation expense for 2019 = $228,900 * 91000/(818000-55000) = $27,300

Book value at the end of 2019 = $228,900 - $27,300 = $201,600

Equipment:

Cost to depreciated = $91,000 - $7,000 = $84,000

Depreciation expense for 2018 = $84,000 * 55000/600000 = $7,700

Book value at the end of 2018 = $91,000 - $7,700 = $83,300

Remaining cost to be depreciated at beginning of 2019 = $83,300 -$7,000 = $76,300

Depreciation expense for 2019 = $76,300 * 91000/(818000-55000) = $9,100

Book value at the end of 2019 = $83,300 - $9,100 = $74,200